OpenStudy (anonymous):

completely factor the expression by grouping: d^2-du-5d+5u

OpenStudy (anonymous):

d^2-du-5d+5u = d[d-u]-5[d-u] = (d-u)(d-5)

OpenStudy (anonymous):

think of d^2 =dxd , 5d= 5xd, du=dxu and 5u=5xu then put them together d^2-du-5d+5u=(d-5)(d-u)

OpenStudy (anonymous):

or d^2-du-5d+5u=d(d-5)-u(d-5)=(d-5)(d-u)

OpenStudy (anonymous):

hope you get a better understanding of how i do it kelly..

OpenStudy (anonymous):

Thanks a bunch!!!

OpenStudy (anonymous):

ywc..have fun..lol